perpendicular to y=2/5x-11 and has a y-intercept of 2
Answer by nerdybill(6948)

For a line to be perpendicular to another, their slopes must be "negative reciprocal".
.
Looking at:
y=(2/5)x-11
It is in the form of "slope-intercept" form of:
y = mx + b
where
m = slope
b = y-intercept
.
Therefore we KNOW that the slope is 2/5.
To find the "negative reciprocal":
(2/5)x = -1
x = -5/2
.
Now, we know the "slope" of the new line AND
the problem said, it has a "y-intercept of 2"
so, plug it all into:
y = mx + b
y = (-5/2)x + 2